Adv. Math. Econ. 10, 31-49 (2007)

Capital-labor substitution and indeterminacy in continuous-time two-sector models

Jean-Philippe Garnier^{1}, Kazuo Nishimura^{2} and Alain Venditti^{3}

1 Université de la Méditerranée, GREQAM, Marseille 13002, France 2 Institute of Economic Research, Kyoto University, Kyoto 606-8501, Japan 3 CNRS, GREQAM, Marseille 13002, France

Received: August 8, 2006 Revised: November 24, 2006 JEL classification: C62, E32, O41 Mathematical Subject Classification (2000): 37C70, 37N40, 37C75, 39A11 Abstract. The aim of this paper is to discuss the role of the elasticity of capitallabor substitution on the local determinacy properties of the steady state in a two-sector economy with CES technologies and sector-specific externalities. Key words: sector-specific externalities, constant returns, capital-labor substitution, indeterminacy